# Exploring Spatial Schema Intuitions in Large Language and Vision Models

### v1

@@ 1 Introduction
  • Crucial differences to human answers remain: the output of small and base models often shows low correlations, LLMs tend to give polarized responses (selecting either 1 or 7 on a scale), and VLMs show lowest correlations with open-source VLMs showing none
@@ 2.1 Image Schemas in Cognition & Language
2.2 Image Schemas in Language Models?
@@ 3 Related Work
3.1 Human Behavior Experiments with LLMs
Using LLMs as human stand-in participants for psychology experiments has recently gained attention Dillion et al. (2023); Harding et al. (2023); Aher et al. (2023). Such use can be motivated by wanting to generate initial hypotheses for an experiment, pilot a new design, and gain insight into human cognition based on the assumption that LLMs trained on a large amount of human-generated text will produce similar output to that of human participants Dillion et al. (2023). For instance, Dillion et al. (2023), who propose such a use, report a high correlation of 0.95 between human answers and GPT-3.5 answers on moral judgment tasks. At the same time, they acknowledge that current LLMs are bad at capturing variation and diversity present in human responses and are biased towards responses of people from certain countries, economic backgrounds, genders, etc. Harding et al. (2023) critique the use of LLMs to replace human participants and question the informativeness of the LLM’s output.
Another motivation to simulate psychological experiments with LLMs is to gain insights not into human cognition but into the capabilities of language models themselves. Reproducing various experiments with LLMs, one can compare the LLM output with how humans behaved in the real experiment, thereby establishing the “human-likeness” of the model’s text generations. The usefulness of such experiments has also been suggested with respect to psycholinguistics, where experiments can show what properties of language can be successfully processed, reproduced, or generated by LLMs Houghton et al. (2023). Dentella et al. (2023), for example, show that LLMs fail at distinguishing grammatical and ungrammatical sentences in a similar way to people. Other studies find that on many psychology tasks, the LLM output is comparable to human answers, even showing similar cognitive biases Hagendorff et al. (2022); Dasgupta et al. (2022). Hagendorff et al. (2022) show that these cognitive biases tend to vanish when experimenting with the most recent models, such as ChatGPT and GPT-4. Aher et al. (2023) extend the idea of repeating prominent experiments with LLMs. Specifically, they not only look at a single output of an LLM given some experiment prompt but try to simulate different demographics by prompting the model multiple times with different personas attached to each prompt. In general, one needs to be aware that the same or similar experiments or tests might have been part of the training data.
@@ 3.2 NLP for Image Schemas
LLMs have only sparingly been used for image schema-related tasks. Initial research on the computational processing of image schemas was restricted to spectral cluster analysis Gromann and Hedblom (2017), whereas more recent work Wachowiak et al. (2022); Wachowiak and Gromann (2022) uses the language model XLM-RoBERTa Conneau et al. (2020). Wachowiak and Gromann (2022) show that language models can be fine-tuned to classify sentences based on eight image schema classes, with an accuracy between 57% and 80% depending on the language. Although their model requires first seeing more than 1,000 correctly annotated samples, these results indicate that, in principle, it is possible for a neural model trained only on text to pick up the pattern indicating a specific image schema in natural language. However, in contrast to the experiments in our work, the language samples they use are manually collated or created by experts. The psycho-linguistic experiments that build the foundation for the present work take a different approach, letting multiple people annotate phrases based on intuitions and felt relatedness. Accordingly, the expert-annotated image schemas are annotated with a single discrete label per sample, while the psycho-linguistic experiments inspiring this study lead to fine-grained annotations using ordinal scales for five different image schemas per phrase. Kamath et al. (2023) test various VLMs for whether they can correctly classify simple spatial configurations, such as A being left, right, under, or over B. Thus, they test for image schematic relations in their original physical form rather than the abstract extensions. They find models to perform poorly on the task, with the best model only achieving an accuracy of 60%, while humans achieve 99%, one of the reasons being that prepositions occur infrequently and in an ambiguous manner in the training data. Similar work by Jassim et al. (2023) shows shortcomings of VLMs’ understanding of spatial configurations and intuitive physics when given different images or video inputs of simulated spatial scenes.
@@ 4 Method
Recreating a prompt that closely mirrors the approach of the original paper would involve consolidating all the stimuli into a single list and instructing the models to associate each stimulus with one of the image schemas. However, all preliminary experiments revealed that this comprehensive prompt yielded impractical and unmeaningful responses. Frequently, models redundantly reproduced identical outputs for each item in the list. Consequently, we opted to refine our approach by providing the models with a single stimulus per input prompt.
@@
To extract ratings from an LLM, we retrieve the most likely answer generated, i.e. the answer received with a temperature of 0 in the OpenAI API or a top_k of 1 with HuggingFace. Alternatively, one could consider the probability for each valid answer, the numbers between 1 and 7, and compute the sum of each number weighted by their likelihood, normalized by the sum of all seven numbers’ likelihoods. While this way takes the LLM’s uncertainty into account, it also requires seven times the amount of compute as simply taking the top answer. Only minimal result variations led us to choose the cost-effective methodology of selecting only the top answer.
@@ 7 Conclusion
This exploratory study is the first of its kind to reproduce psycholinguistic experiments in order to explore spatial schema intuitions. Moreover, it provides evidence that LLMs are able to reflect those intuitions in different tasks and setups. The results also point out that despite the duo-modality, VLMs do not encode the spatial understanding as effectively as their textual origins. Future research aims to assess models using novel, collected stimuli, ensuring no prior exposure during training. Our current findings provide a foundation for formulating precise, testable hypotheses in subsequent experiments. Additionally, we would like to extend this line of research to include a multilingual analysis.
@@ 8 Limitations
Additionally, our reliance on psycholinguistic data introduces notable limitations. Firstly, the temporal aspect of the data is a concern, given its age range of 30 to 23 years. Language evolves over time, and the potential disparities between our data and contemporary linguistic trends may affect the applicability of our results. Therefore, we plan to replicate the studies to gauge temporal robustness.
@@ 9 Considerations and Impact of the Work
Overall, text generations with the OpenAI API cost 25.85\$. Text generations with LLaMA models were run on two clusters. In those experiments, we utilized 8 NVIDIA RTX A6000 (48GB) GPUs for a 4-hour runtime, each with a power consumption of 300 W, resulting in an estimated total power consumption of 9.6 kWh and a CO2 emission of approximately 3.984 kg. Additionally, we used 2 NVIDIA A100 (40GB) GPUs for a 39-hour runtime, also consuming 300 W each, contributing to a combined total power consumption of 33 kWh and a CO2 emission of approximately 13.7 kg.
@@
Balance
Consider the notion of Balance. Balance refers to your sense of symmetry or stability relative to some point within your body.
Verticality
Consider the notion of Verticality. Verticality refers to the sense of an extension along an up–down orientation.
Center–Periphery
Consider the notion of Center–Periphery. Center–periphery refers to the experience of some objects or events as central while surrounding objects and events are peripheral or to the outside.
Resistance
Consider the notion of Resistance. Resistance refers to the experience of your body opposing some external force.
Linkage
Consider the notion of Linkage. Linkage refers to the perception of a connection between objects or events.
@@
Support
In the case of the use of “on” in “the book is on the desk”: the Support relation refers to the desk supporting the book.
Pressure
In the case of the use of “on” in “the book is on the desk”: the Pressure relation refers to the book exerting some pressure on the desk.
Constraint
In the case of the use of “on” in “the book is on the desk”: the Constraint relation refers to the desk constraining the possible motions of the book.
Covering
In the case of the use of “on” in “the book is on the desk”: the Covering relation refers to the book concealing the part of the desk that is under the book.
Visibility
In the case of the use of “on” in “the book is on the desk”: the Visibility relation refers to the book being visible on the desk.
@@
Textual
Given the concepts: [DIRECTION WORDS]. For the concept that best represents the event '[ACTION WORD]', I would choose '
Pseudo-Visual
Given the event '[ACTION WORD]', which of the following arrows best represents this event: [UNICODE ARROWS]. A research participant would choose the arrow:
Visual
[Instruction: You are shown four images. Select one of the four images to answer the question.]
Image [IMAGE LABEL]:[IMAGE]
Image [IMAGE LABEL]:[IMAGE]
Image [IMAGE LABEL]:[IMAGE]
Image [IMAGE LABEL]:[IMAGE]
Question: Which is the best image to describe the event '[ACTION WORD]'? Answer: For '[ACTION WORD]' it is Image
GPT-4chat
[TEXTUAL or PSEUDO-VISUAL PROMPT] You are a participant in a research experiment. Even if the answer is subjective, provide it. Do not say it is subjective.

### v2

@@ 1 Introduction
  • Crucial differences to human answers remain: the output of small and base models often shows low correlations, LLMs tend to give polarized responses (selecting either 1 or 7 on a scale), and VLMs show lowest correlations with open-source VLMs showing none.
@@ 2.1 Image Schemas in Cognition & Language
2.2 Image Schemas in Language Models
@@ 3 Related Work
3.1 Human Behavioral Experiments with Large Language Models
Using LLMs as human stand-in participants for psychology experiments has recently gained attention Futrell et al. (2019); Linzen and Baroni (2021); Dillion et al. (2023); Harding et al. (2023); Aher et al. (2023). Such use can be motivated by wanting to generate initial hypotheses for an experiment, pilot a new design, and gain insight into human cognition based on the assumption that LLMs trained on a large amount of human-generated text will produce similar output to that of human participants Dillion et al. (2023). For instance, Dillion et al. (2023), who propose such a use, report a high correlation of 0.95 between human answers and GPT-3.5 answers on moral judgment tasks. At the same time, they acknowledge that current LLMs are bad at capturing variation and diversity present in human responses and are biased towards responses of people from certain countries, economic backgrounds, genders, etc. Harding et al. (2023) critique the use of LLMs to replace human participants and question the informativeness of the LLM’s output.
Another motivation to simulate psychological experiments with LLMs is to gain insights not into human cognition but into the capabilities of language models themselves. Reproducing various experiments with LLMs, one can compare the LLM output with how humans behaved in the real experiment, thereby establishing the “human-likeness” of the model’s text generations. The usefulness of such experiments has also been suggested with respect to psycholinguistics, where experiments can show what properties of language can be successfully processed, reproduced, or generated by LLMs Houghton et al. (2023).
A survey by Linzen and Baroni (2021) presents studies of neural networks’ syntactic abilities and their broader implications for linguistic theory. Dentella et al. (2023), for example, show that LLMs fail at distinguishing grammatical and ungrammatical sentences in a similar way to people. Futrell et al. (2019) test four neural network language models on artificial sentences with syntactically complex structures (subordinate clauses and the Garden Path effect) to analyze their syntactic representations. Their findings indicate that LSTMs trained on large datasets represent syntactic states comparably to an RNNG trained on a small dataset, while an LSTM trained on a small dataset performs poorly or only weakly. Other studies find that on many psychology tasks, the LLM output is comparable to human answers, even showing similar cognitive biases Hagendorff et al. (2022); Dasgupta et al. (2022). Hagendorff et al. (2022) show that these cognitive biases tend to vanish when experimenting with the most recent models, such as ChatGPT and GPT-4. Aher et al. (2023) extend the idea of repeating prominent experiments with LLMs. Specifically, they not only look at a single output of an LLM given some experiment prompt but try to simulate different demographics by prompting the model multiple times with different personas attached to each prompt. Generally, one needs to be aware that the same or similar experiments might have been in the training data.
@@ 3.2 NLP for Image Schemas
LLMs have only sparingly been used for image schema-related tasks. Initial research on the computational processing of image schemas was restricted to spectral cluster analysis Gromann and Hedblom (2017), whereas more recent work Wachowiak et al. (2022); Wachowiak and Gromann (2022) uses the language model XLM-RoBERTa Conneau et al. (2020). Wachowiak and Gromann (2022) show that language models can be fine-tuned to classify sentences based on eight image schema classes, with an accuracy between 57% and 80% depending on the language of the data. Although their model requires first seeing more than 1,000 correctly annotated samples, these results indicate that, in principle, it is possible for a neural model trained only on text to pick up the pattern indicating a specific image schema in natural language. However, in contrast to the experiments in our work, the language samples they use are manually collated or created by experts. The psycho-linguistic experiments that build the foundation for the present work take a different approach, letting multiple people annotate phrases based on intuitions and felt relatedness. Accordingly, the expert-annotated image schemas are annotated with a single discrete label per sample, while the psycho-linguistic experiments inspiring this study lead to fine-grained annotations using ordinal scales for five different image schemas per phrase. Kamath et al. (2023) test various VLMs for whether they can correctly classify simple spatial configurations, such as A being left, right, under, or over B. Thus, they test for image schematic relations in their original physical form rather than the abstract extensions. They find models to perform poorly on the task, with the best model only achieving an accuracy of 60%, while humans achieve 99%, one of the reasons being that prepositions occur infrequently and in an ambiguous manner in the training data. Similar work by Jassim et al. (2023) shows shortcomings of VLMs’ understanding of spatial configurations and intuitive physics when given different visual inputs of simulated spatial scenes.
From the corpus of related works delving into psycholinguistic studies probing spatial schema intuitions, we selected three experiments for comparison, each exhibiting variations in word class and context. Experiment 1 Gibbs et al. (1994) showcases a verb (stand) within context, Experiment 2 Beitel et al. (2001) features a preposition (on) within context, and Experiment 3 Richardson et al. (2001) employs different verbs without context (see Fig. 1). Our rationale for excluding other studies stemmed from their failure to offer sufficient variation or to provide access to their original questionnaire. Moreover, the chosen experiments span a spectrum of distinct spatial schemas.
@@ 4 Method
Recreating a prompt that closely mirrors the approach of the original paper would involve consolidating all the stimuli into a single list and instructing the models to rate each stimulus in combination with each image schema. We tried conditions in which the model had to rate all stimuli for all image schemas and all stimuli for a single schema. Additionally, we tried averaging multiple of those runs based on different stimuli orders. However, all preliminary experiments revealed that this comprehensive prompt yielded impractical or even unmeaningful responses. Frequently, models redundantly reproduced identical outputs for each item in the list. Consequently, we opted to refine our approach by providing the models with a single stimulus per input prompt.
@@
To extract ratings from an LLM, we retrieve the most likely answer generated, i.e. the answer received with a temperature of 0 in the OpenAI API or a top_k of 1 with HuggingFace. Alternatively, one could consider the probability for each valid answer, the numbers between 1 and 7, and compute the sum of each number weighted by their likelihood, normalized by the sum of all seven numbers’ likelihoods. While this takes the LLM’s uncertainty into account, it also requires seven times the amount of compute compared to simply taking the top answer. As results only varied minimally, we chose the cost-effective methodology of selecting only the top answer.
@@ 5 Results
Despite GPT-4 showing the highest correlations, it still generates some answers strongly deviating from human spatial intuitions. For example, GPT-4 fails to relate a sense of VERTICALITY to the phrase “the barometer stands at 30 inches” (scoring it 1 compared to the participant average of 4.71) or BALANCE to the phrase “the clock stands on the mantle” (1 compared to 4.46). Generally, it can be observed that many of the outliers are caused when the model gives too low ratings, i.e., 1 or 2.
@@ 7 Conclusion
This exploratory study is the first of its kind to reproduce psycholinguistic experiments in order to explore spatial schema intuitions. Moreover, it provides evidence that LLMs are able to reflect those intuitions in different tasks and setups. The results also point out that despite the duo-modality, VLMs do not encode the spatial understanding as effectively as their textual origins. Future research aims to assess models using novel, collected stimuli, ensuring no prior exposure during training. Our current findings provide a foundation for formulating precise, testable hypotheses in subsequent experiments. Additionally, we would like to extend this line of research to include a multilingual analysis. In conclusion, our study not only sheds light on the disparities between LLMs and human cognition but also paves the way for new research perspectives in understanding and refining language models. By providing empirical evidence of these disparities, we advocate for a deeper exploration into the limitations of current models and the development of novel approaches to bridge the gap between artificial intelligence and human intelligence.
@@ 8 Limitations
Furthermore, the original experiments have a limited demographic, which we hereby report: Gibbs et al. (1994) with 27 undergraduate students, U.S. university, native English speakers. Beitel et al. (2001) with 24 undergraduate students, U.S. university, native English speakers. Richardson et al. (2001) with 173 undergraduate students, U.S. university, no further information is provided.
Additionally, our reliance on psycholinguistic data introduces notable limitations. Firstly, the temporal aspect of the data is a concern, given its age range of 30 to 23 years. Language evolves over time, and the potential disparities between our data and contemporary linguistic trends may affect the applicability of our results. Therefore, we plan to replicate the studies to gauge temporal robustness. Moreover, we acknowledge that the reproduced studies solely feature the English language and multilingual analysis is subject of future work.
@@ 9 Considerations and Impact of the Work
Overall, text generations with the OpenAI API cost 25.85\$. Text generations with LLaMA models were run on two clusters. In those experiments, we utilized 8 NVIDIA RTX A6000 (48GB) GPUs for a 4-hour runtime, each with a power consumption of 300 W, resulting in an estimated total power consumption of 9.6 kWh and a CO2 emission of approximately 3.984 kg. Additionally, we used 2 NVIDIA A100 (40GB) GPUs for a 39-hour runtime, also consuming 300 W each, contributing to a combined total power consumption of 33 kWh and a CO2 emission of ˜13.7 kg.
Acknowledgements
Lennart Wachowiak was supported by UK Research and Innovation (EP/S023356/1), in the UKRI Centre for Doctoral Training in Safe and Trusted Artificial Intelligence, as well as the King’s Institute for Artificial Intelligence.
@@ Acknowledgements
Balance Consider the notion of Balance. Balance refers to your sense of symmetry or stability relative to some point within your body.
Verticality Consider the notion of Verticality. Verticality refers to the sense of an extension along an up–down orientation.
Center–Periphery Consider the notion of Center–Periphery. Center–periphery refers to the experience of some objects or events as central while surrounding objects and events are peripheral or to the outside.
Resistance Consider the notion of Resistance. Resistance refers to the experience of your body opposing some external force.
Linkage Consider the notion of Linkage. Linkage refers to the perception of a connection between objects or events.
@@
Support In the case of the use of “on” in “the book is on the desk”: the Support relation refers to the desk supporting the book.
Pressure In the case of the use of “on” in “the book is on the desk”: the Pressure relation refers to the book exerting some pressure on the desk.
Constraint In the case of the use of “on” in “the book is on the desk”: the Constraint relation refers to the desk constraining the possible motions of the book.
Covering In the case of the use of “on” in “the book is on the desk”: the Covering relation refers to the book concealing the part of the desk that is under the book.
Visibility In the case of the use of “on” in “the book is on the desk”: the Visibility relation refers to the book being visible on the desk.
@@
Table A4: List of action words from Richardson et al. (2001). These words are used for Experiment 3.
@@
Textual Given the concepts: [DIRECTION WORDS]. For the concept that best represents the event '[ACTION WORD]', I would choose '
Pseudo-Visual Given the event '[ACTION WORD]', which of the following arrows best represents this event: [UNICODE ARROWS]. A research participant would choose the arrow:
Visual [Instruction: You are shown four images. Select one of the four images to answer the question.] Image [IMAGE LABEL]:[IMAGE] Image [IMAGE LABEL]:[IMAGE] Image [IMAGE LABEL]:[IMAGE] Image [IMAGE LABEL]:[IMAGE] Question: Which is the best image to describe the event '[ACTION WORD]'? Answer: For '[ACTION WORD]' it is Image
GPT-4chat [TEXTUAL or PSEUDO-VISUAL PROMPT] You are a participant in a research experiment. Even if the answer is subjective, provide it. Do not say it is subjective.